# Lorentz transformation in Quantum Mechanics

### v2

@@ Abstract.
The compatibility of special relativity and Quantum Mechanics has been questioned by several authors. The origin of this tension can be traced back mainly to the introduction of the measurement processes and the corresponding wave function reduction, which play a crucial role in Quantum Mechanics. We approach this problem within a recent proposal for a model of Quantum Mechanics, where the measurement is explicitly described as a specific stochastic process. This implements ordinary Quantum Mechanics, where measurement and reduction are treated as phenomenological events of unknown origin without any physical justification. To state clearly the question, we first discuss and establish the effect of a Lorentz transformation on a generic wave function in space-time. This allows to formulate the relativistic version of the model. We then consider few thought experiments in order to analyze to what extent Quantum Mechanics follows relativistic invariance and find the specific critical points where non invariance possibly occurs. The relevance of this analysis on the experimental observations is also discussed.
@@
Quantum Mechanics (QM) and special relativity are the basis of the theoretical description of all physical phenomena, with the exception of the gravitational ones. For a coherent theory of natural phenomena it is essential that they are compatible. The well established relativistic Quantum Field Theory (QFT) [1] is indeed a coherent formulation of QM in the relativistic framework. However, when the wave functions are introduced to represent physical systems, evidences of possible incompatibility clearly arise, in particular when measurements are considered in space-time. The most known example is the spin correlation measurement between entangled particles, the so called EPR paradox [2, 3]. The quantum mechanical calculation of the correlation perfectly agrees with the experimental finding, but it is performed assuming instantaneous reduction of the wave function, which amounts to an instantaneous ”propagation” of the correlation. Of course a correlation is not a signal, since the result of the spin measurement is stochastic and it cannot be controlled to send a message. However it looks difficult to accommodate this instantaneous correlation within a relativistic formalism. In fact, as noticed in ref. [4], the order in which one of the particle affects the state of the other depends on the reference frame if Lorentz transformation (LT) are used, which hints a violation of the ”cause - effect” temporal order. As it is well known [5, 6] the correlation itself is not relativistic invariant since it depends on the reference frame, but the violation of Bell’s inequality and of the local realism is still valid. This and other examples will be discussed in section 5.
@@
In this paper we approach the problem within the model of ref. [7] for the completion of QM. The model is devised to describe the physical processes that are kept outside the standard formalism, in particular the strong measurements. The basic assumption is the introduction for each wave function of a nonstandard component, in the sense of nonstandard analysis, and of a nonstandard space-time lattice which is the support of the wave function. Under a set of additional assumptions, that are coherent with the basic one, it turns out that a Ito stochastic process is automatically present in the standard sector. This stochastic dynamics has to be added to the standard linear evolution ruled by Schrödinger equation. It can trigger the reduction of the wave function in agreement with Born’s rule, which is a process not contained in the formalism of standard QM. The introduction of the lattice and the corresponding stochastic dynamics is the basic element of the model, since no standard external perturbation can produce the reduction of a wave function [8].
@@
In section 2 we sketch the relativistic extension of the model. In secs. 3,4 we establish in general the LT of a wave function on the basis of QFT. In sec. 5 we study few cases where measurements in space-time play an essential role and we discuss to what extent and in which sense QM and relativity are compatible.
@@
At the classical (non-quantal) level it would be difficult to question LT, therefore we will assume that they are valid in the standard sector, i.e. for the standard part of the space-time coordinates. On the other hand we cannot use LT for the nonstandard part, because this would mix the order of the infinitesimal steps in the space-time lattice. This requires a more refined interpretation of the meaning of the space-time lattice. First of all it has to be stressed that the LT of Eq. (2) include an implicit assumption, often omitted. As explicitly mentioned in the original paper by A. Einstein [9], it is supposed that identical copies of a basic ruler and a basic clock are provided to each reference frame, that are used to measure the space-time coordinates. The values of the space-time coordinates of an event are in fact obtained by taking the ratio of the intervals along the axis from the origin and the basic space-time units. This is a essential assumption, otherwise it would have no meaning to relate the coordinates of two different reference frames. This is equivalent to consider these basic elements not affected by the LT. To illustrate this point, one can refer to the example of the previous section about the contraction of length, where the measure of a stick length in the moving frame is obtained from the interval between the stick extremities taken at the same time. The interval has to be compared with the unit of length, which of course is not taken as the LT of the one of the rest frame, but it is just the one belonging to the moving frame.
It is natural to identify the space-time infinitesimal steps as the basic ruler and basic ”clock tick”, which are therefore independent of the reference frame. In this way in different reference frames an event has different space-time standard coordinates, in agreement with LT, but the same infinitesimal steps for the nonstandard lattice. This interpretation does not violate Lorentz invariance, since it is in accordance with the interpretation of the standard LT. This identification is equivalent to consider the infinitesimal steps as the minimal space-time intervals one can consider in the space-time manifold. Being infinitesimal, they are of course not directly measurable, but they can affect specific physical phenomena, as in the non-relativistic case. Because the finite part of the nonstandard straight line (or any curve) is the disjoint union of all the monadΠ of the real axis, each coordinate can be seen as the juxtaposition of an unlimited number of these basic elements along the space-time axis. All the juxtapositions which end in a given monadΠ give the same standard value for the coordinate.
@@
which means that in the new reference frame the wave function is the same original function in the new coordinates. It has to be stressed that the time $t\,^{\prime}$ is different for each $\mathbf{x}$. In QFT for a field extended in a finite space-time region this non-uniqueness of the time under LT is an unavoidable consequence. It means simply that the field at a given time in a reference frame is related by a LT to the field taken at different times in another reference frame. However, if $\phi(x)$ is meant to represent the wave function of a particle in a given reference frame, the effect of a LT to another reference frame poses a straightforward question. In fact, in QM the wave function at a given time is assumed to represent the quantum state of the system at that definite time. The non-uniqueness of the time implies the impossibility of connecting by a LT a definite quantum state in two different reference frames. Indeed, in general no correspondence can be established between the quantum state represented by the wave function $\phi(t,\mathbf{x})$ and the quantum state represented by a wave function $\phi(t\,^{\prime},\mathbf{x\,^{\prime}})$ at a given $t\,^{\prime}$. A correspondence can be present only between the whole evolution of the wave function as observed in the two reference frames. The main reason is that the wave packet is an extended object in 3-D space and evolves in both reference frames, which marks a difference with the case of a rigid body at rest in one of the two reference frames. A further obvious difference is the impossibility to ”observe” a wave function like a physical object. In some limiting case the correspondence between the states in the two reference frames can be established, at least approximately. Suppose that the wave packet is ”almost” two dimensional, i.e. in one direction the size of the wave packet can be neglected. Then in the reference frame where the x-axis is along this special direction all the points of the wave packet have the same x-coordinate and a boost along x axis will bring to a wave packet with only one time coordinate, i.e. a legitimate quantum state. Similarly, it can happen that in definite physical conditions the overall size of the wave packet is small enough that the spread of the time values under the LT of Eq. (4) can be neglected. This is the case of e.g. the semi-classical approximation.
@@
Let us consider the elementary process of detection of a particle by an apparatus which is at rest in the considered reference frame, e.g. a measurement of position of a particle described by a wave packet with a finite space extension. The measurement will produce the reduction of the wave function, e.g. the localization of the wave function in a restricted space region. As seen by another reference frame, which is moving with respect to that reference frame, it will be described by performing a LT on the space-time coordinates. Then, as discussed in the previous section 4, we have to apply the transformation to the whole evolution of the wave function. Therefore the LT will involve the wave function both before and after the reduction time in the original reference frame. As a consequence the reduction process as seen by the moving frame will not be anymore instantaneous, but it could be described in general as a gradual transformation of the incoming wave packet into the reduced form. The reference frame where the experimental apparatus is at rest plays therefore a special role. In this sense the reduction process is not completely relativistic invariant [4]. This conclusion does not depend on the particular model used for the reduction process, provided it is instantaneous in the special reference frame, but it is just a consequence of the spatial extension of the wave function.
@@
A LT could reverse the order of the possible detections, in which case detector 2 will fire, while detector 1 will be silent. In more details, let us consider first the reference frame where the detectors are at rest and arrange the detectors at equal distances from the source. Then the two particles will arrive at the corresponding detectors simultaneously. Let us first switch off the detectors, so we can consider the two arrivals without the occurrence of detection. If we now perform a LT along the x-axis, i.e we observe the system from a moving frame, the two arrivals will be not anymore simultaneous, but they will be observed with a time delay $\Delta t=\gamma\beta\Delta x$, where $\gamma=1/\sqrt{1-\beta^{2}}$ and $\Delta x$ is the distance along the x-axis of the two detectors in the original standing frame. Then, according to the sign of the velocity $\beta$ between the two frames, the arrivals will be in different temporal order. Switching on the detectors, this means that only one of the detectors will be operating.
Along these lines one can consider the EPR [2] experiments, which had numerous realizations and improvements in order to finally circumvent all possible pitfalls and loopholes. An extensive list of references on the experiments can be found in ref. [11]. Of course the main interest on these experiments is the observed violation of Bell’s inequality, which excludes the possibility of a local realism in QM. The origin of the correlations involved in the Bell’ s inequality is essentially threefold, the reduction of the two particle wave function under measurement, the extension of the wave function and its integrity. The QM calculation of the correlations assumes an instantaneous reduction and it is in perfect agreement with observations. This gives support to the use of the wave function in the QM formalism, which then acquires some degree of realism. This result do not depend on possible model for the reduction process, it must be only compatible with the instantaneity assumption, like the model of ref. [7].
@@ 6 Summary and conclusions
In the realistic interpretations of Quantum Mechanics one introduces the wave function as a fundamental tool at the basis of its formalism and of its practical applications. Historically the wave function was devised in order to explain the observations of diffraction and interference of elementary particles, in particular electrons. Since then it became the universal mathematical object in the formalism to be used to interpret and design all the experiments where Quantum Mechanics is involved. In the applications a special role is played by the measurement processes, where the so called wave function reduction occurs. The reduction is not part of the mathematical formalism but introduced as a separate postulate without an explicit description. Recently a model was proposed [7] that extends the standard Quantum Mechanics to incorporate in the formalism the reduction process. In order to devise the relativistic version of the model, we first establish the effect of a LT on a generic wave function. On this basis we discuss the LT of the reduction process and we consider several thought experiments, to enlight possible incompatibility between relativity and Quantum Mechanics. One of the main feature of the LT on a wave function is the impossibility of establishing a direct relation between the observations of a quantum state in two different reference frames. Only the overall evolution of a wave function in two reference frames can be connected by a LT. The reason for this peculiarity is just the unavoidable changes under LT of the time interval between two non coincident space-time events. As a consequence, in the case of a generic detection, the reduction time, which is vanishing in the detector reference frame, becomes finite under a LT. The length of this time depends on the size of the wave function undergoing the reduction.
Analyzing some thought experiments some apparent paradoxical situations seem to occur. In particular, in a two-particle systems, the arrangement of two detectors can be such that only one particle is detected, but which one of the two depends on the reference frame. In other words, the only detector that is actually firing depends on the reference frame. Despite this can appear paradoxical, it does not violate any basic principle. We finally considered the EPR experiment, and it is found that it is necessary to keep the reduction process as distinct from the other physical processes, i.e. as belonging to a separate category. Despite that, if a model, like the one proposed in ref. [7], is able to explain the reduction of a general wave function in a measuring process, there is no logical or factual paradox. However, one is forced to accept that the correlations established in measurements on an entangled system cannot be explained by a fully relativistic theory. This conclusion points out an existing tension between Quantum Mechanics and relativity.

### v3

@@ Abstract.
The compatibility of special relativity and Quantum Mechanics has been questioned by several authors. The origin of this tension can be traced back mainly to the introduction of the measurement processes and the corresponding wave function reduction, which play a crucial role in Quantum Mechanics. We approach this problem with the help of a recent proposal for a model of Quantum Mechanics, where the measurement is explicitly described as a specific stochastic process. This implements ordinary Quantum Mechanics, where measurement and reduction are treated as phenomenological events of unknown origin without any physical justification. To state clearly the question in general, we first discuss and establish the effect of a Lorentz transformation on a generic wave function in space-time. Alongside the analysis we formulate the relativistic version of the model. We then consider few thought experiments in order to analyze to what extent Quantum Mechanics follows relativistic invariance and find the specific critical points where non invariance possibly occurs. The analysis can shade light to the interpretation of the existing experimental observations.
@@
Quantum Mechanics (QM) and special relativity are the basis of the theoretical description of all physical phenomena, with the exception of the gravitational ones. For a coherent theory of natural phenomena it is essential that they are compatible. The well established relativistic Quantum Field Theory (QFT) [1] is indeed a coherent formulation of QM in the relativistic framework. However, when the wave functions are introduced to represent physical systems, evidences of possible incompatibility clearly arise, in particular when measurements are considered in space-time. Since the laws of ordinary Quantum Mechanics (QM) cannot be questioned, because they are in perfect agreement with experiments, it is of great interest to clarify to what extent the laws of special relativity is compatible with them. This is surely not a new subject, but there are points that are missing in the literature, which are worth to be considered and discussed. The paper does not present new theory or numerical predictions or to ”solve” the paradoxes, since they are independent of any model. The most known example is the spin correlation measurement between entangled particles, the so called EPR paradox [2, 3]. The quantum mechanical calculation of the correlation perfectly agrees with the experimental finding, but it is performed assuming instantaneous reduction of the wave function, which amounts to an instantaneous ”propagation” of the correlation. Of course a correlation is not a signal, since the result of the spin measurement is stochastic and it cannot be controlled to send a message. However it looks difficult to accommodate this instantaneous correlation within a relativistic formalism. In fact, as noticed in ref. [4], the order in which one of the particle affects the state of the other depends on the reference frame if Lorentz transformation (LT) are used, which hints a violation of the ”cause - effect” temporal order. As it is well known [5, 6] the correlation itself is not relativistic invariant since it depends on the reference frame, but the violation of Bell’s inequality and of the local realism is still valid. This and some thought experiments will be discussed in section 5.
@@
In this paper we approach the problem with the help of the model of ref. [7] for the completion of QM. The model is devised to describe the physical processes that are kept outside the standard formalism, in particular the strong measurements. The basic assumption is the introduction for each wave function of a nonstandard component, in the sense of nonstandard analysis, and of a nonstandard space-time lattice which is the support of the wave function. The lattice has infinitesimal steps, where the time step is of higher order than the space step. Under a set of additional assumptions, that are coherent with the basic one, it turns out that a particular Ito stochastic process is automatically present in the standard sector. This stochastic dynamics has to be added to the standard linear evolution ruled by Schrödinger equation. The corresponding stochastic differential equations is then obtained under the only condition of the conservation of the norm. This equation is able to produce both the wave function reduction process and the corresponding Born’ s rule for the probability, which in ordinary QM are without explanations. The introduction of the lattice and the corresponding stochastic dynamics is the basic element of the model, since no standard external perturbation can produce the reduction of a wave function [8].
Possible predictions of the model, that could give evidence for the presence of the nonstandard lattice, are briefly discussed in refs. [7, 9].
@@
The paper is organized as follows. In section 2 we sketch the relativistic extension of the model. In secs. 3,4 we establish in general the LT of a wave function on the basis of QFT. In sec. 5 we study few cases where measurements in space-time play an essential role and we discuss to what extent and in which sense QM and relativity are compatible.
@@
At the classical (non-quantal) level it would be difficult to question LT, therefore we will assume that they are valid in the standard sector, i.e. for the standard part of the space-time coordinates. On the other hand we cannot use LT for the nonstandard part, because this would mix the order of the infinitesimal steps in the space-time lattice. In fact LT involves both space and time intervals, and the infinitesimal of higher order for the time step is vanishing small with respect to the infinitesimal for the space step. As can be seen from the LT equations (2) the transformed time step would become of the same order of the space step, and the model would be contradictory. This requires a more refined interpretation of the meaning of the space-time lattice. First of all it has to be stressed that the LT of Eq. (2) include an implicit assumption, often omitted. As explicitly mentioned in the original paper by A. Einstein [10], it is supposed that identical copies of a basic ruler and a basic clock are provided to each reference frame, that are used to measure the space-time coordinates. The values of the space-time coordinates of an event are in fact obtained by taking the ratio of the intervals along the axis from the origin and the basic space-time units. This is a essential assumption, otherwise it would have no meaning to relate the coordinates of two different reference frames. This is equivalent to consider these basic elements not affected by the LT. To illustrate this point, one can refer to the example of the previous section about the contraction of length, where the measure of a stick length in the moving frame is obtained from the interval between the stick extremities taken at the same time. The interval has to be compared with the unit of length, which of course is not taken as the LT of the one of the rest frame, but it is just the one belonging to the moving frame.
It is natural to identify the space-time infinitesimal steps as the basic ruler and basic ”clock tick”, which are therefore independent of the reference frame. In this way in different reference frames an event has different space-time standard coordinates, in agreement with LT, but the same infinitesimal steps for the nonstandard lattice. This interpretation does not violate Lorentz invariance, since it is in accordance with the interpretation of the standard LT. This identification is equivalent to consider the infinitesimal steps as the minimal space-time intervals one can consider in the space-time manifold. As explained above, this interpretation of the space-time steps is the only one that preserves the model, any other choice would lead to inconsistent results. Being infinitesimal, they are of course not directly measurable, but they can affect specific physical phenomena, as in the non-relativistic case. Because the finite part of the nonstandard straight line (or any curve) is the disjoint union of all the monadΠ of the real axis, each coordinate can be seen as the juxtaposition of an unlimited number of these basic elements along the space-time axis. All the juxtapositions which end in a given monadΠ give the same standard value for the coordinate.
@@
which means that in the new reference frame the wave function is the same original function in the new coordinates. It has to be stressed that the time $t\,^{\prime}$ is different for each $\mathbf{x}$. In QFT for a field extended in a finite space-time region this non-uniqueness of the time under LT is an unavoidable consequence. It means simply that the field at a given time in a reference frame is related by a LT to the field taken at different times in another reference frame. However, if $\phi(x)$ is meant to represent the wave function of a particle in a given reference frame, the effect of a LT to another reference frame poses a straightforward question. In fact, in QM the wave function at a given time is assumed to represent the quantum state of the system at that definite time. The non-uniqueness of the time implies the impossibility of connecting by a LT a definite quantum state in two different reference frames. Indeed, in general no correspondence can be established between the quantum state represented by the wave function $\phi(t,\mathbf{x})$ and the quantum state represented by a wave function $\phi(t\,^{\prime},\mathbf{x\,^{\prime}})$ at a given $t\,^{\prime}$. A correspondence can be present only between the whole evolution of the wave function as observed in the two reference frames. The main reason is that the wave packet is an extended object in 3-D space and evolves in both reference frames, which marks a difference with the case of a rigid body at rest in one of the two reference frames. A further obvious difference is the impossibility to ”observe” a wave function like a physical object. This is one of the main conclusions of the paper, which is not explicitly present in the literature. It is a straightforward consequence of how a LT acts on a wave function and the interpretation of the wave function in QM as representing a quantum state. Stated in a different way, if one performs a LT on a wave function from a reference frame, in the new space-time coordinates the wave functions at different times appear, and no quantum state can be defined. Despite that, the range of time involved in the set of transformed wave functions can be easily estimated. Let us consider a wave function in a given reference frame $\phi(t,x)$ at a given time $t$ which has a space extension of the order $\Delta x$. A LT to a reference frame moving with velocity $\beta$ will introduce a set of wave functions spanning a time interval of the order $\Delta t\,=\,\beta\gamma\Delta x$, where $\gamma=1/\sqrt{1-\beta^{2}}$.
In some limiting case the correspondence between the states in the two reference frames can be established, at least approximately. Suppose that the wave packet is ”almost” two dimensional, i.e. in one direction the size of the wave packet can be neglected. Then in the reference frame where the x-axis is along this special direction all the points of the wave packet have the same x-coordinate and a boost along x axis will bring to a wave packet with only one time coordinate, i.e. a legitimate quantum state. Similarly, it can happen that in definite physical conditions the overall size of the wave packet is small enough that the spread of the time values under the LT of Eq. (4) can be neglected. This is the case of e.g. the semi-classical approximation.
@@
Let us consider the elementary process of detection of a particle by an apparatus which is at rest in the considered reference frame, e.g. a measurement of position of a particle described by a wave packet with a finite space extension. The measurement will produce the reduction of the wave function, e.g. the localization of the wave function in a restricted space region. As seen by another reference frame, which is moving with respect to that reference frame, it will be described by performing a LT on the space-time coordinates. Then, as discussed in the previous section 4, we have to apply the transformation to the whole evolution of the wave function. Therefore the LT will involve the wave function both before and after the reduction time in the original reference frame. As a consequence the reduction process as seen by the moving frame will not be anymore instantaneous, but it could be described in general as a set of wave functions that include a gradual transformation of the incoming wave packet into the reduced form. However the new description of the process as a set wave functions cannot be related to a quantum state evolution, since no quantum state can be indeed identified. On the other way around, if we consider the wave function as seen by the moving frame at a given time, by LT from the fixed reference frame it will receive contributions from both reduced and non-reduced wave functions. Of course this change of description occurs for any other process that can happen to the original wave function. The reference frame where the experimental apparatus is at rest plays therefore a special role. In this sense the reduction process is not completely relativistic invariant [4]. This conclusion does not depend on the particular model used for the reduction process, provided it is instantaneous in the special reference frame, but it is just a consequence of the spatial extension of the wave function.
@@
A LT could reverse the order of the possible detections, in which case detector 2 will fire, while detector 1 will be silent. In more details, let us consider first the reference frame where the detectors are at rest and arrange the detectors at equal distances from the source. Then the two particles will arrive at the corresponding detectors simultaneously. Let us first switch off the detectors, so we can consider the two arrivals without the occurrence of detection. If we now perform a LT along the x-axis, i.e we observe the system from a moving frame, the two arrivals will be not anymore simultaneous, but they will be observed with a time delay $\Delta t=\gamma\beta\Delta x$, where $\Delta x$ is the distance along the x-axis of the two detectors in the original standing frame. Then, according to the sign of the velocity $\beta$ between the two frames, the arrivals will be in different temporal order. Switching on the detectors, this means that only one of the detectors will be operating.
Along these lines one can consider the EPR [2] experiments, which had numerous realizations and improvements in order to finally circumvent all possible pitfalls and loopholes. An extensive list of references on the experiments can be found in ref. [12]. Of course the main interest on these experiments is the observed violation of Bell’s inequality, which excludes the possibility of a local realism in QM. The origin of the correlations involved in the Bell’ s inequality is essentially threefold, the reduction of the two particle wave function under measurement, the extension of the wave function and its integrity. The measurement selects one two-particles spin (or polarization) state between the two that appears in the entangled superposition. The effect of LT on the values of the spin correlations where considered extensively in ref. [5, 6]. Here we focus on the space-time correlations of the measurements by the two opposite detectors, not considered in [5, 6]. The QM calculation of the correlations assumes an instantaneous reduction and it is in perfect agreement with observations, for all the detector distances considered in the experiments, in particular when the detectors are at equal distance from the source. This gives support to the use of the wave function in the QM formalism, which then acquires some degree of realism. This result do not depend on possible model for the reduction process, it must be only compatible with the instantaneity assumption, like the model of ref. [7]. Let me stress that instantaneity property is not a trivial requirement, since, to the author knowledge, hardly other models are able to ’explain’ it.
@@ 6 Summary and conclusions
In the realistic interpretations of Quantum Mechanics one introduces the wave function as a fundamental tool at the basis of its formalism and of its practical applications. Historically the wave function was devised in order to explain the observations of diffraction and interference of elementary particles, in particular electrons. Since then it became the universal mathematical object in the formalism to be used to interpret and design all the experiments where Quantum Mechanics is involved. In the applications a special role is played by the measurement processes, where the so called wave function reduction occurs. The reduction is not part of the mathematical formalism but introduced as a separate postulate without an explicit description. Recently a model was proposed [7] that extends the standard Quantum Mechanics to incorporate in the formalism the reduction process. In order to devise the relativistic version of the model, we first establish the effect of a LT on a generic wave function. On this basis we discuss the LT of the reduction process and we consider several thought experiments, to enlight possible incompatibility between relativity and Quantum Mechanics. One of the main feature of the LT on a wave function is the impossibility of establishing a direct relation between the observations of a quantum state in two different reference frames. Only the overall evolution of a wave function in two reference frames can be connected by a LT. The reason for this peculiarity is just the unavoidable changes under LT of the time interval between two non coincident space-time events. As a consequence, in the case of a generic detection, the reduction time, which is vanishing in the detector reference frame, appears finite under a LT. The length of this time depends on the size of the wave function undergoing the reduction.
Analyzing some thought experiments some apparent paradoxical situations seem to occur. In particular, in a two-particle systems, the arrangement of two detectors can be such that only one particle is detected, but which one of the two depends on the reference frame. In other words, the only detector that is actually firing depends on the reference frame. Despite this can appear paradoxical, it does not violate any basic principle. We finally considered the EPR experiment, and it is pointed out that the corresponding reduction process is not relativistic invariant, and therefore a special role is played by the reference frame where the experimental apparatus is at rest and where the probabilities, according to the Born’ s rule, have to be calculated. Despite that, if a model, like the one proposed in ref. [7], is able to explain the reduction of a general wave function in a measuring process, there is no logical or factual paradox. These conclusions specify to what extent Quantum Mechanics and relativity are compatible.